# Defending Jailbreak Attack in VLMs via Cross-modality Information Detector

### v2

@@ Abstract
Vision Language Models (VLMs) extend the capacity of LLMs to comprehensively understand vision information, achieving remarkable performance in many vision-centric tasks. Despite that, recent studies have shown that these models are susceptible to jailbreak attacks, which refer to an exploitative technique where malicious users can break the safety alignment of the target model and generate misleading and harmful answers. This potential threat is caused by both the inherent vulnerabilities of LLM and the larger attack scope introduced by vision input. To enhance the security of VLMs against jailbreak attacks, researchers have developed various defense techniques. However, these methods either require modifications to the model’s internal structure or demand significant computational resources during the inference phase. Multimodal information is a double-edged sword. While it increases the risk of attacks, it also provides additional data that can enhance safeguards. Inspired by this, we propose Cross-modality Information DEtectoR (CIDER), a plug-and-play jailbreaking detector designed to identify maliciously perturbed image inputs, utilizing the cross-modal similarity between harmful queries and adversarial images. This simple yet effective cross-modality information detector, CIDER, is independent of the target VLMs and requires less computation cost. Extensive experimental results demonstrate the effectiveness and efficiency of CIDER, as well as its transferability to both white-box and black-box VLMs.
Defending Jailbreak Attack in VLMs via Cross-modality Information Detector
Yue Xu1thanks: equal contribution , Xiuyuan Qi1* , Zhan Qin2 , Wenjie Wang1 thanks: W.Wang is the corresponding author. 1School of Information Science and Technology, ShanghaiTech University 2College of Computer Science and Technology, Zhejiang University {xuyue2022,qixy1,wangwj1}@shanghaitech.edu.cn, qinzhan@zju.edu.cn
@@ 1 Introduction
The remarkable advancements in Large Language Models (LLMs) have significantly improved performance benchmarks in various natural language processing (NLP) tasks Achiam et al. (2023); Touvron et al. (2023); Zhao et al. (2023); Chiang et al. (2023). To extend the capacities and open up the potentials of LLMs in comprehensively understanding diverse types of data, such as visual information, researchers have developed Vision Language Models (VLMs) that integrate visual modalities to handle vision-centric tasks. VLMs use LLMs as a core, complemented by modal-specific encoders and projectors, enabling them to process, reason, and generate outputs from multimodal data Yin et al. (2023); Dai et al. (2024); Bai et al. (2023). A typical VLM architecture is illustrated in Figure 1.
Figure 1: The illustration of a typical VLM architecture.
@@
The widespread adoption of VLMs in various applications brings significant safety challenges, particularly due to inherited vulnerabilities from traditional LLMs, such as the susceptibility to jailbreak attacks Carlini et al. (2024); Li et al. (2024); Qi et al. (2024). Jailbreak attacks refer to an exploitative technique where malicious users can craft sophisticated-designed prompts to lead LLMs to answer misleading or harmful questions, effectively breaking the alignment and bypassing the model’s safeguard. Various jailbreak attack algorithms targeting LLMs have been proposed, which can be categorized into template-based Deng et al. (2024); Chao et al. (2023); Li et al. (2023) and optimize-based Zou et al. (2023) approaches.
@@
Additionally, VLMs not only inherit the vulnerabilities of LLMs but also become more susceptible to jailbreak attacks due to their integration of the visual modality. On the one hand, jailbreak attacks against VLMs can originate from both the textual and visual modalities, significantly broadening the scope of potential adversarial examples Gong et al. (2023); Shayegani et al. (2023). On the other hand, recent research indicates that fine-tuning VLMs to learn the vision modality can cause LLMs to disregard their previously learned safety alignment Zong et al. (2024).
@@
Figure 2: The workflow of safeguarding VLM against jailbreak attacks via CIDER.
@@
The existing jailbreak attacks on VLMs can be categorized into two strategies. One is white-box optimization-based attacks, which define a loss function to generate imperceptible perturbations in the image modality Carlini et al. (2024); Qi et al. (2024); Niu et al. (2024). The other is black-box strategies including typographically transforming harmful queries into images such as FigStep Gong et al. (2023) or adding related images containing harmful text such as QR Liu et al. (2023).
@@
From the defense perspective, optical character recognition (OCR) can serve as an effective detection tool for the second strategy but fails when defending against optimization-based adversarial attacks. In addition, Zong et al. (2024) creates a vision-language dataset named VLGuard containing both safe and unsafe queries and images, which can be used to fine-tune VLMs for improved safety against jailbreak attacks. However, the effectiveness of VLGuard is only tested on FigStep attack and it requires the model to be white-box to fine-tune. Zhang et al. (2023) proposed a mutation-based jailbreaking detection framework named Jailguard. However, the performance of Jailguard heavily relies on the VLMs’ original safety alignment, and it significantly increases computational costs during the inference phase.
@@
Multimodal information is a double-edged sword: while it increases the risk of attacks, it also provides additional data that helps enhance safeguards. Inspired by this potential, we propose Cross-modality Information DEtectoR (CIDER), a plug-and-play jailbreaking detector designed to identify maliciously perturbed image inputs, specifically targeting optimization-based jailbreak attacks that are more imperceptible and susceptible. The intuition is that optimization-based perturbations break the VLM’s safeguards by capturing the main harmful content in the malicious query. As a result, the semantic distance between a harmful query and an adversarially perturbed image is significantly smaller than that between a harmful query and a clean image.
@@
Directly utilizing the difference between clean and adversarial images on the semantic distance to harmful query is challenging, as the absolute value of the distance varies across different harmful queries. To address this issue, we incorporate a denoiser to preprocess the vision modality, using the relative shift in the semantic distance before and after denoising to reflect the difference between clean and adversarial images. As shown in Figure 2, the key insight of CIDER is to identify whether an image is adversarially perturbed based on the semantic similarity between image and text modality before and after denoising ($\langle\bm{E_{\textit{text}}},\bm{E_{\textit{img(o)}}}\rangle-\langle\bm{E_{\textit{text}}},\bm{E_{\textit{img(d)}}}\rangle$). If the image modality is not perturbed, the semantic similarity between text and image remains stable. However, the adversarially perturbed image designed for jailbreak will experience a significant drop. By setting a threshold based on this change, we can effectively detect adversarially perturbed images aimed at jailbreaking VLMs. The detailed intuition is elaborated in Section 2.
@@
As a pre-detection module encapsulated before any VLMs, the key advantage of CIDER is its plug-and-play nature, making it independent of the target model. Additionally, timely inference is crucial for safeguarding VLMs. CIDER achieves this by adding only denoiser procedures, ensuring efficient without introducing significant inference latency.
@@
In this work, we propose CIDER, an effective and efficient pre-detection module that denoises and inspects each input image. For images identified as adversarially perturbed for jailbreak purposes (where the semantic shift exceeds a pre-defined threshold), the VLM will refuse to generate a response. Images deemed normal will be processed along with the text input for model inference by the VLM. The workflow of safeguarding VLMs against jailbreak attacks using CIDER is illustrated in Figure 2. Our contribution can be summarized as follows:
@@
  • We propose a plug-and-play jailbreaking detector, CIDER, which can effectively safeguard VLMs while incurring almost no additional computational overhead.
  • Extensive experiments validate that CIDER outperforms the baseline method, achieving a higher detection success rate while reducing the computational cost as well. Experimental results also demonstrate its strong transferability across both white-box and black-box VLMs and attack methods.
@@ 2 Intuition: Cross-modality information is a double-edged sword
While multimodal information aggravates model vulnerability to jailbreak attacks, it also provides additional information for defense. The design of CIDER is based on the intuition that optimization-based jailbreak attacks break the VLM’s safeguards by sharing harmful content in the malicious query to the image modality. Consequently, the adversarially perturbed image is closer to the harmful query in the semantic space than the clean images. To support this intuition, we first explain the fundamentals of the optimization-based jailbreak attacks on VLMs. Then, we design a few experiments to explore how cross-modal analysis can help safeguard VLMs, and we analyze the semantic difference between clean and adversarial images relative to harmful queries, both before and after denoising.
2.1 Preliminaries: Optimization-based Jailbreak Attacks on VLMs
Optimization-based VLM jailbreaking is similar to adversarial attacks on image classification tasks Goodfellow et al. (2014), with the primary difference being the difference in the loss function. Specifically, given a dataset $D=\{(q,a)\}$ where $q$ represents the harmful queries and $a$ is the corresponding targeted answers, the attacker aims to find an adversarial image $x_{adv}$ that can encourage the VLM $\mathcal{F}$ to generate $a$ when inputting $q$ along with $x_{adv}$. The objective can be formulated as:
@@ 2.1 Preliminaries: Optimization-based Jailbreak Attacks on VLMs
where $\mathcal{F}(a|q,x_{adv})$ represents the likelihood that the VLM $\mathcal{F}$ generate answer $a$ when given the adversarial image $x_{adv}$ and the query $q$.
@@ 2.2 Experimental Setup
We design a series of experiments to explore how cross-modality information can help safeguard VLMs and to analyze the semantic difference between clean and adversarial images to harmful queries, before and after denoising. We utilize the image and text encoder of the state-of-the-art VLM LLaVA-v1.5-7B Liu et al. (2024) to capture the semantic meanings. To measure the semantic similarity, we employed cosine similarity which is a standard metric widely used in information retrieval and natural language processing Park et al. (2020); Pal et al. (2021). In terms of denoiser, we incorporate a diffusion-based denoiser Nichol and Dhariwal (2021) to preprocess the image modality.
The inputs to the VLMs consist of two modalities: images and text queries. For malicious queries, we utilize the validation set proposed in the Harmbench framework Mazeika et al. (2024), which contains 40 textual harmful behaviors across 7 semantic categories. For images, we use 5 adversarial images generated by an optimization-based jailbreak attack Qi et al. (2024) and 5 clean images from ImageNet Deng et al. (2009). As a result, we have 200 adversarial text-image pairs and 200 clean pairs.
@@ 2.3 Findings
Subsequently, we applied denoising to each image 350 times, assessing cosine similarity with harmful queries every 50 iterations (visualization of the denoising is relegated to Appendix A). Figure 3(c) illustrates how cosine similarity between harmful query and adversarial images decreases as denoising progresses, indicating a reduction in harmful information. Despite this reduction, denoised adversarial images, when tested with harmful text inputs in the VLM, still enabled a significant number of queries to jailbreak. Thus, while denoising mitigates harmful information in images, it does not eliminate their adversarial properties.
@@ 3.1 Overview
CIDER is implemented on top of the VLMs to defense optimization-based adversarial jailbreak attacks. Figure 2 presents the overview of the CIDER pipeline. Specifically, given a text-image input pair, denoted as <text, img(o)>, CIDER calculates the embeddings of text and image modalities, denoted as $\bm{E_{\textit{text}}}$ and $\bm{E_{\textit{img(o)}}}$. Then, the built-in denoiser in CIDER will perform 350 denoising iterations on the image(o), calculating the denoised embeddings every 50 iterations, denoted as $\mathcal{E}=\bm{E_{\textit{img(d)}}}$. The img(o) will be identified as an adversarial example if any $\bm{E_{\textit{img(d)}}}\in\mathcal{E}$ satisfy the following condition:
@@
Input: $img(o)$: input image, $text$: input query, $\mathcal{F}$: target VLM, $\tau$: predefined threshold.
@@ 4 Experiment
In this section, we begin by outlining the configurations of our experiments, including the models, datasets, baselines, and evaluation metrics. We then evaluate the effectiveness and efficiency of CIDER, comparing with the baseline methods. Next, we discuss the trade-off between robustness and utility. Finally, we demonstrate the generalization of our method.
@@ 4.1 Configurations
Models. Note that CIDER is an auxiliary model that is independent to the VLMs. We use LLaVA to capture the semantic meaning of each modality, but CIDER can be plugged into any other VLMs. To demonstrate the effectiveness of CIDER, we test the detection and defense performance on four open-source VLMs, including LLaVA-v1.5-7B Liu et al. (2024), MiniGPT4 Zhu et al. (2023), InstructBLIP Dai et al. (2024), and Qwen-VL Bai et al. (2023), as well as the API-access VLM, GPT4V Achiam et al. (2023).
Figure 5: ASR of base VLM, defending with CIDER and defending with Jailguard
Datasets. Similar to the dataset used in Section 2.2, we generate 800 adversarial text-image pairs utilizing the 160 harmful queries in Harmbench Mazeika et al. (2024) and adversarial images provided by Qi et al. (2024). To further demonstrate CIDER will not destroy the original utilities on the normal queries, we also evaluate the utility of CIDER protected VLMs on MM-Vet benchmark Yu et al. (2023), which examines 6 core vision language capabilities, including recognition, optical character recognition (OCR), knowledge, language generation, spatial awareness, and math.
Baseline and evaluation metrics. We use Jailguard Zhang et al. (2023) as a baseline, which is a SoTA mutation-based jailbreak detection strategy that protects the VLMs at the inference stage. We involve four evaluation metrics to demonstrate the performance of defending methods from different aspects. From the perspective of the effectiveness of CIDER, we incorporate detection success rate (DSR) and Attack success rate (ASR). DSR represents the proportion of adversarial examples $\mathcal{D}$ that can be successfully detected:
@@
ASR is a standard evaluation metric indicating the proportion of samples that can successfully jailbreak VLM $\mathcal{F}$ and generate harmful contents, which can be stated as:
@@
$\mathcal{G}$ refers to an LLM classifier Mazeika et al. (2024) that determines the harmfulness of a response. $\mathbb{I}_{harm}$ and $\mathbb{I}_{harm}$ represent the adversarial and harmful indicator. In terms of efficiency, we measure the time cost of processing a single query. In addition, to evaluate the model utility on regular tasks, responses, we incorporate an online evaluator MM-Vet-Evaluator (2024) provided along with MM-Vet benchmark, which utilizes GPT-4 to generate a soft grading score from 0 to 1 for each answer.

### v4

# Cross-modality Information Check for Detecting Jailbreaking in Multimodal Large Language Models

@@ Abstract
Multimodal Large Language Models (MLLMs) extend the capacity of LLMs to understand multimodal information comprehensively, achieving remarkable performance in many vision-centric tasks. Despite that, recent studies have shown that these models are susceptible to jailbreak attacks, which refer to an exploitative technique where malicious users can break the safety alignment of the target model and generate misleading and harmful answers. This potential threat is caused by both the inherent vulnerabilities of LLM and the larger attack scope introduced by vision input. To enhance the security of MLLMs against jailbreak attacks, researchers have developed various defense techniques. However, these methods either require modifications to the model’s internal structure or demand significant computational resources during the inference phase. Multimodal information is a double-edged sword. While it increases the risk of attacks, it also provides additional data that can enhance safeguards. Inspired by this, we propose Cross-modality Information DEtectoR (CIDER), a plug-and-play jailbreaking detector designed to identify maliciously perturbed image inputs, utilizing the cross-modal similarity between harmful queries and adversarial images. CIDER is independent of the target MLLMs and requires less computation cost. Extensive experimental results demonstrate the effectiveness and efficiency of CIDER, as well as its transferability to both white-box and black-box MLLMs. The resource is available at https://github.com/PandragonXIII/CIDER.
Cross-modality Information Check for Detecting Jailbreaking in Multimodal Large Language Models
Yue Xu1thanks: equal contribution , Xiuyuan Qi1* , Zhan Qin2 , Wenjie Wang1 thanks: W.Wang is the corresponding author. 1School of Information Science and Technology, ShanghaiTech University 2The State Key Laboratory of Blockchain and Data Security, Zhejiang University {xuyue2022,qixy1,wangwj1}@shanghaitech.edu.cn, qinzhan@zju.edu.cn
@@ 1 Introduction
The remarkable advancements in Large Language Models (LLMs) have significantly improved performance benchmarks in various natural language processing (NLP) tasks Achiam et al. (2023); Touvron et al. (2023); Zhao et al. (2023); Chiang et al. (2023). To extend the capacities and open up the potentials of LLMs in comprehensively understanding diverse types of data, such as visual information, researchers have developed Multimodal Large Language Models (MLLMs) that integrate visual modalities to handle vision-centric tasks. MLLMs use LLMs as a core, complemented by modal-specific encoders and projectors, enabling them to process, reason, and generate outputs from multimodal data Yin et al. (2023); Dai et al. (2024); Bai et al. (2023). A typical MLLM architecture is illustrated in Figure 1.
Figure 1: The architecture of a typical MLLM.
@@
The widespread adoption of MLLMs in various applications brings significant safety challenges, particularly due to inherited vulnerabilities from traditional LLMs, such as the susceptibility to jailbreak attacks Carlini et al. (2024); Li et al. (2024); Qi et al. (2024). Jailbreak attacks refer to an exploitative technique where malicious users can craft sophisticated-designed prompts to lead LLMs to answer misleading or harmful questions, effectively breaking the safety alignment. Various jailbreak attack algorithms targeting LLMs have been proposed, which can mainly be categorized into template-based Deng et al. (2024); Chao et al. (2023); Li et al. (2023) and optimization-based Zou et al. (2023) approaches.
@@
Additionally, MLLMs not only inherit the vulnerabilities of LLMs but also become more susceptible to jailbreak attacks due to their integration with other modalities. On the one hand, jailbreak attacks against MLLMs can originate from both the textual and visual modalities, significantly broadening the scope of potential adversarial examples Gong et al. (2023); Shayegani et al. (2023). On the other hand, recent research indicates that fine-tuning MLLMs to learn the vision modality can cause LLMs to disregard their previously learned safety alignment Zong et al. (2024).
@@
Figure 2: The workflow of safeguarding MLLM against jailbreak attacks via CIDER.
@@
The existing jailbreak attacks on MLLMs can be categorized into two strategies. One is white-box optimization-based attacks, which define a loss function to generate imperceptible perturbations in the image modality Carlini et al. (2024); Qi et al. (2024); Niu et al. (2024). The other is black-box strategies including typographically transforming harmful queries into images such as FigStep Gong et al. (2023) or adding related images containing harmful text such as QR Liu et al. (2023).
@@
From the defense perspective, optical character recognition (OCR) can serve as an effective detection tool for the second strategy but fails when defending against optimization-based adversarial attacks. In addition, Zong et al. (2024) creates a vision-language dataset named VLGuard containing both safe and unsafe queries and images, which can be used to fine-tune MLLMs for improved safety against jailbreak attacks. However, the effectiveness of VLGuard is only tested on FigStep attack and it requires the model to be white-box to fine-tune. Zhang et al. (2023) proposed a mutation-based jailbreaking detection framework named Jailguard. However, the performance of Jailguard heavily relies on the MLLMs’ original safety alignment, and it significantly increases computational costs during the inference phase.
@@
Multimodal information is a double-edged sword: while it increases the risk of attacks, it also provides additional data that helps enhance safeguards. Inspired by this potential, we propose Cross-modality Information DEtectoR (CIDER), a plug-and-play jailbreaking detector designed to identify maliciously perturbed image inputs, specifically targeting optimization-based jailbreak attacks that are more imperceptible and susceptible. The intuition is that optimization-based perturbations break the MLLM’s safeguards by capturing the main harmful content in the malicious query. As a result, the semantic distance between a harmful query and an adversarially perturbed image is significantly smaller than that between a harmful query and a clean image.
@@
Directly utilizing the difference between clean and adversarial images on the semantic distance to harmful query is challenging, as the absolute value of the distance varies across different harmful queries. To address this issue, we incorporate a denoiser to preprocess the vision modality, using the relative shift in the semantic distance before and after denoising to reflect the difference between clean and adversarial images. As shown in Figure 2, the key insight of CIDER is to identify whether an image is adversarially perturbed based on the semantic similarity between image and text modality before and after denoising ($\langle\bm{E_{\textit{text}}},\bm{E_{\textit{img(o)}}}\rangle-\langle\bm{E_{\textit{text}}},\bm{E_{\textit{img(d)}}}\rangle$). If the image modality is not perturbed, the semantic similarity between text and image remains stable. However, the adversarially perturbed image designed for jailbreak will experience a significant drop. By setting a threshold based on this change, we can effectively detect adversarially perturbed images aimed at jailbreaking MLLMs. The detailed intuition is elaborated in Section 2.
@@
As a pre-detection module encapsulated before any MLLMs, the key advantage of CIDER is its plug-and-play nature, making it independent of the target model. Additionally, timely inference is crucial for safeguarding MLLMs. CIDER achieves this with denoising procedures, ensuring efficiency and minimal latency.
@@
In this work, we propose CIDER, an effective and efficient pre-detection module that denoises and inspects each input image. For images identified as adversarially perturbed for jailbreak purposes (where the semantic shift exceeds a pre-defined threshold), the MLLM will refuse to generate a response. Images deemed normal will be processed along with the text input for model inference by the MLLM. The workflow of safeguarding MLLMs against jailbreak attacks using CIDER is illustrated in Figure 2. Our contribution can be summarized as follows:
@@
  • We propose a plug-and-play jailbreaking detector, CIDER, which can effectively safeguard MLLMs while incurring almost no additional computational overhead.
  • Extensive experiments validate that CIDER outperforms the baseline method, achieving a higher detection success rate while reducing the computational cost as well. Experimental results also demonstrate its strong transferability across both white-box and black-box MLLMs and attack methods.
@@ 2 Intuition: Cross-modality information is a double-edged sword
While multimodal information aggravates model vulnerability to jailbreak attacks, it also provides additional information for defense. The design of CIDER is based on the intuition that optimization-based jailbreak attacks break the MLLM’s safeguards by sharing harmful content in the malicious query to the image modality. Consequently, the adversarially perturbed image is closer to the harmful query in the semantic space than the clean images. To support this intuition, we first explain the fundamentals of the optimization-based jailbreak attacks on MLLMs. Then, we design a few experiments to explore how cross-modal analysis can help safeguard MLLMs, and we analyze the semantic difference between clean and adversarial images relative to harmful queries, both before and after denoising.
2.1 Preliminaries: Optimization-based Jailbreak Attacks on MLLMs
Optimization-based MLLM jailbreaking is similar to adversarial attacks on image classification tasks Goodfellow et al. (2014), with the primary difference being the difference in the loss function. Specifically, given a dataset $D=\{(q,a)\}$ where $q$ represents the harmful queries and $a$ is the corresponding targeted answers, the attacker aims to find an adversarial image $x_{adv}$ that can encourage the MLLM $\mathcal{F}$ to generate $a$ when inputting $q$ along with $x_{adv}$. The objective can be formulated as:
@@ 2.1 Preliminaries: Optimization-based Jailbreak Attacks on MLLMs
where $\mathcal{F}(a|q,x_{adv})$ represents the likelihood that the MLLM $\mathcal{F}$ generate answer $a$ when given the adversarial image $x_{adv}$ and the query $q$.
@@ 2.2 Experimental Setup
We design a series of experiments to explore how cross-modality information can help safeguard MLLMs and to analyze the semantic difference between clean and adversarial images to harmful queries, before and after denoising. We utilize the image and text encoder of the state-of-the-art MLLM LLaVA-v1.5-7B Liu et al. (2024) to capture the semantic meanings. To measure the semantic similarity, we employed cosine similarity which is a standard metric widely used in information retrieval and natural language processing Park et al. (2020); Pal et al. (2021). In terms of denoiser, we incorporate a diffusion-based denoiser Nichol and Dhariwal (2021) to preprocess the image modality.
The inputs to the MLLMs consist of two modalities: images and text queries. For malicious queries, we utilize the validation set proposed in the Harmbench framework Mazeika et al. (2024), which contains 40 textual harmful behaviors across 7 semantic categories. For images, we use 5 adversarial images generated by an optimization-based jailbreak attack Qi et al. (2024) and 5 clean images from ImageNet Deng et al. (2009). As a result, we have 200 adversarial text-image pairs and 200 clean pairs.
@@ 2.3 Findings
Subsequently, we applied denoising to each image 350 times, assessing cosine similarity with harmful queries every 50 iterations (visualization of the denoising is relegated to Appendix A). Figure 3(c) illustrates how cosine similarity between harmful query and adversarial images decreases as denoising progresses, indicating a reduction in harmful information. Despite this reduction, denoised adversarial images, when tested with harmful text inputs in the MLLM, still enabled a significant number of queries to jailbreak. Thus, while denoising mitigates harmful information in images, it does not eliminate their adversarial properties.
@@ 3.1 Overview
CIDER is implemented on top of the MLLMs to defense optimization-based adversarial jailbreak attacks. Figure 2 presents the overview of the CIDER pipeline. Specifically, given a text-image input pair, denoted as <text, img(o)>, CIDER calculates the embeddings of text and image modalities, denoted as $\bm{E_{\textit{text}}}$ and $\bm{E_{\textit{img(o)}}}$. Then, the built-in denoiser in CIDER will perform 350 denoising iterations on the image(o), calculating the denoised embeddings every 50 iterations, denoted as $\mathcal{E}=\bm{E_{\textit{img(d)}}}$. The img(o) will be identified as an adversarial example if any $\bm{E_{\textit{img(d)}}}\in\mathcal{E}$ satisfy the following condition:
@@
Input: $img(o)$: input image, $text$: input query, $\mathcal{F}$: target MLLM, $\tau$: predefined threshold.
@@ 4 Experiment
In this section, we begin by outlining the configurations of our experiments, including the models, datasets, baselines, and evaluation metrics. We then evaluate the effectiveness and efficiency of CIDER, comparing with the baseline methods. Next, we discuss the trade-off between robustness and utility, and the choice of denoising method. Finally, we demonstrate the generalization of our method.
@@ 4.1 Configurations
Models. Note that CIDER is an auxiliary model that is independent of the MLLMs. We use LLaVA to capture the semantic meaning of each modality, but CIDER can be plugged into any other MLLMs. To demonstrate the effectiveness of CIDER, we test the detection and defense performance on four open-source MLLMs, including LLaVA-v1.5-7B Liu et al. (2024), MiniGPT4 Zhu et al. (2023), InstructBLIP Dai et al. (2024), and Qwen-VL Bai et al. (2023), as well as the API-access MLLM, GPT4V Achiam et al. (2023).
Figure 5: ASR of base MLLM, defending with CIDER and defending with Jailguard
Datasets. Similar to the dataset used in Section 2.2, we generate 800 adversarial text-image pairs utilizing the 160 harmful queries in Harmbench Mazeika et al. (2024) and adversarial images provided by Qi et al. (2024). To further demonstrate CIDER will not destroy the original utilities on the normal queries, we also evaluate the utility of CIDER protected MLLMs on MM-Vet benchmark Yu et al. (2023), which examines 6 core vision language capabilities, including recognition, optical character recognition (OCR), knowledge, language generation, spatial awareness, and math.
Baseline and evaluation metrics. We use Jailguard Zhang et al. (2023) as a baseline, which is a SoTA mutation-based jailbreak detection strategy that protects the MLLMs at the inference stage. We involve four evaluation metrics to demonstrate the performance of defending methods from different aspects. From the perspective of the effectiveness of CIDER, we incorporate detection success rate (DSR) and Attack success rate (ASR). DSR represents the proportion of adversarial examples $\mathcal{D}$ that can be successfully detected:
@@
ASR is a standard evaluation metric indicating the proportion of samples that can successfully jailbreak MLLM $\mathcal{F}$ and generate harmful contents, which can be stated as:
@@
$\mathcal{G}$ refers to an LLM classifier Mazeika et al. (2024) that determines the harmfulness of a response. $\mathbb{I}_{adv}$ and $\mathbb{I}_{harm}$ represent the adversarial and harmful indicator. In terms of efficiency, we measure the time cost of processing a single query. In addition, to evaluate the model utility on regular tasks, and responses, we incorporate an online evaluator MM-Vet-Evaluator (2024) provided along with MM-Vet benchmark, which utilizes GPT-4 to generate a soft grading score from 0 to 1 for each answer.